If a number is multiplied by 5 the result is 2 less than 3 times the sum of the number and 6. Find the number.
x=8 but I get 2.
Here's how I set up the equation:
5x = (3x-2)+6

Answer by macston(5194)

5x+2=3(x+6)
5x+2=3x+18
2x=16
x=8
CHECK:
5x+2=3(x+6)
5(8)+2=3(8+6)
42=3(14)
42=42